OpenStudy (tiffany_rhodes):

Find the integral of 1/(x^2-a^2) dx, where a is not equal to 0. Use partial fraction decomposition. I started the problem, but I'm getting nowhere.

OpenStudy (anonymous):

wat u got so far?

OpenStudy (anonymous):

did u factor the denominator?

OpenStudy (tiffany_rhodes):

Yes, I factored the denominator using the difference of squares. The factors would be (x+a)(x-a)

OpenStudy (anonymous):

ok... so \(\large \frac{1}{(x+a)(x-a)}=\frac{A}{x+a} +\frac{B}{x-a} \) now you need to get rewrite the right side so the fractions "can" be combined (common denominators)..

OpenStudy (anonymous):

what equation do you get as a result?

OpenStudy (anonymous):

\(\large \frac{1}{(x+a)(x-a)}=\frac{A}{x+a} +\frac{B}{x-a} \) clearing the fractions, we have: \(\large \frac{1}{(x+a)(x-a)}=\frac{A(x-a)}{(x+a)(x-a)} +\frac{B(x+a)}{(x-a)(x+a)} \) \(\large 1=A(x-a)+B(x+a) \) ok so far?

OpenStudy (tiffany_rhodes):

Yes, I've got everything so far.

OpenStudy (anonymous):

now just let x=a, solve for B; then let x=-a, solve for A.... what u got A = ??? , B = ???

OpenStudy (tiffany_rhodes):

That's where I got confused. Why do we let x = a?

OpenStudy (anonymous):

to eliminate A so we can solve for B... and we choose specific values of x to make the job easier... notice if we let x=a, then that will eliminate A(x-a) because A(a-a) = 0. so we end up with the equation: \(\large 1=A(a-a)+B(a+a) \) \(\large 1=B(2a) \) ok so far?

OpenStudy (tiffany_rhodes):

Oh okay, yes I got it.

OpenStudy (anonymous):

so, find A and B so we can do the integral... A = B =

OpenStudy (tiffany_rhodes):

A= -1/2a and B= 1/2a

OpenStudy (anonymous):

great... so we have... (hang on)

OpenStudy (anonymous):

\(\large \frac{1}{(x+a)(x-a)}=\frac{A}{x+a} +\frac{B}{x-a} \) \(\large \frac{1}{(x+a)(x-a)}=\frac{(-1/2a)}{x+a} +\frac{(1/2a)}{x-a}=\frac{-1}{2a(x+a)} +\frac{1}{2a(x-a)}=\frac{1}{2a(x-a)}-\frac{1}{2a(x+a)} \) ok???

OpenStudy (tiffany_rhodes):

okay, got it.

OpenStudy (anonymous):

i think u should be able to to the rest (integral) from here....

OpenStudy (tiffany_rhodes):

yes, thank you so much!

OpenStudy (anonymous):

yw.... :)

OpenStudy (tiffany_rhodes):

:)